OpenStudy (anj123):

How many one-to-one functions are there from a set with m elements to one with n elements? I do not understand the solution Solution: First note that when m>n there are no one-to-one functions from a set with m elements to a set with n elements...? What does this part of the solution mean? Why does is matter the value of the elements within the set m and n as long as it is all different numbers?

ganeshie8 (ganeshie8):

Hi

ganeshie8 (ganeshie8):

Consider the situation of 10 boys and 9 girls.

ganeshie8 (ganeshie8):

Is it possible for every boy to have exactly one girl friend ?

OpenStudy (usukidoll):

no cuz one girl will be left out

OpenStudy (usukidoll):

whoops I meant one boy will be left out

ganeshie8 (ganeshie8):

That means we cannot have one to one relationship if we have boys more than girls

ganeshie8 (ganeshie8):

Notice that we don't care if girls are left out here. Only the domain, boys, must all be paired up. There can be some girls in the codomain that don't have a preimage

OpenStudy (anj123):

So were assuming that one set is larger than the other?

OpenStudy (anj123):

It is example 7

OpenStudy (anj123):

okay in the question, it says the value of the function at a(k) can be chosen in n-k+1 ways. I understand the n-k but why did they add 1?

terenzreignz (terenzreignz):

are you still in a pinch? ^_^

OpenStudy (sshayer):

\[\left( n-0 \right)\left( n-1 \right)\left( n-2 \right)......\left\{ n-\left( m-1 \right) \right\}=n \left( n-1 \right)\left( n-2 \right)...\left( n-m+1 \right)\]